OpenStudy (chosenmatt):

A school counselor surveyed 90 randomly selected students about the languages they speak. Of the students surveyed, 16 speak more than one language fluently. Based on these results, how many of the 1,800 students at the school can be expected to speak more than one language fluently? A. 113 B. 288 C. 389 D. 320

OpenStudy (chosenmatt):

@iGreen @AriPotta

OpenStudy (chosenmatt):

@perl

OpenStudy (chosenmatt):

@Abhisar

OpenStudy (chosenmatt):

@Nnesha

OpenStudy (chosenmatt):

@Secret-Ninja

OpenStudy (aripotta):

make a proportion :)\[\frac{ part }{ whole }\frac{ 16 }{ 90 }=\frac{ x }{ 1800 }\]cross-multiply and solve for x

OpenStudy (chosenmatt):

im still confused

OpenStudy (aripotta):

do you know how to cross-multiply?

OpenStudy (anonymous):

Cross multiply. \(16 \times 1800\) = \(90 \times x\)

OpenStudy (anonymous):

Basically, just multiply 16 * 1800, then divide that by 90.

OpenStudy (chosenmatt):

oh okay lol

OpenStudy (anonymous):

:P

OpenStudy (chosenmatt):

hi @across

OpenStudy (across):

Did you fully understand what iGreen said?

OpenStudy (chosenmatt):

yes i do thanks!!